I have a final today and there will be similar questions as this one. Can someone please help me understand this problem.
The fracture strength of a certain type of manufactured glass is normally distributed with a mean of 579 MPa with a standard deviation of 14 MPa. (a) What is the probability that a randomly chosen sample of glass will break at less than 579 MPa? (b) More than 590 MPa? (c) Less than 600 MPa?
Thank you

Answer by stanbon(75887)

The fracture strength of a certain type of manufactured glass is normally distributed with a mean of 579 MPa with a standard deviation of 14 MPa. (a) What is the probability that a randomly chosen sample of glass will break at less than 579 MPa? (b) More than 590 MPa? (c) Less than 600 MPa?
---------------------
You have to convert to z-scores.
--------------
(a) What is the probability that a randomly chosen sample of glass will break at less than 579 MPa?
z(579)= (579-579)/14 = 0
P(x < 579) = P(z < 0) = 0.5000
-----------------------------------
(b) More than 590 MPa?
z(590) = (590-579)/14 = 11/14= 0.7857...
P(x > 590) = P(z > 0.7857..) = 0.2160...
-----------------------------------
(c) Less than 600 MPa?
z(600) = (600-579)/14 = 21/14 = 3/2
P(x < 600) = P(z< 3/2) = 0.9332
====================================
Cheers,
Stan H.